your walk to your morning physics lecture takes 10 minutes and as you walk you do mechanical work at a rate of 33W with an efficiency of 6%. How much energy stored in your body ahve you used up over the course of your walk?
1
Expert's answer
2020-04-21T19:13:01-0400

Solution.

"t = 10 minutes;

N = 33W;

η = 6\%;"

The efficiency ratio is the ratio of useful work to full work multiplied by 100%:

"\eta=\dfrac{A_u}{A_f}\sdot100\%;"

"A_f = N \u22c5 t; t =600 seconds ;\n\nA_f =33 \u22c5 600 = 19800J;"

"W = A_u =\dfrac{\eta\sdot A_f }{100\%};W =\dfrac{6\%\sdot 19800J }{100\%}=1188J;"

Answer:"W=1188J."